# : Generalizing Large Language Models for E-commerce

### v1

@@ 6.5 Impact of Training Data Size on
We investigate how the training data size impacts the performance of models. As presented in SectionΒ 3.4, has more than 92 thousand (92K) training samples. In this experiment, we generate 3 smaller training sets of 1K, 10K, and 47K samples. The 1K and 10K samples are generated by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are constructed by randomly selecting 5K samples from each task except for , for which we include all its 2K training samples. The complete results are presented in AppendixΒ H.
@@ 8 Impact Statements
B Instructions
- B.1 Instructions on Attribute Value Extraction ()
- B.2 Instructions on Product Relation Prediction ()
- B.3 Instructions on Product Matching ()
- B.4 Instructions on Sentiment Analysis ()
- B.5 Instructions on Sequential Recommendation ()
- B.6 Instructions on Multi-class Product Classification ()
- B.7 Instructions on Product Substitute Identification ()
- B.8 Instructions on Query Product Ranking ()
- B.9 Instructions on Answerability Prediction ()
- B.10 Instructions on Answer Generation ()
- C Training and Evaluation Details
- D Prompt Templates
- E Analysis on Base Models
- F Complete Experimental Results
- G OOD Evaluation
- H Complete Results for the Analysis on Training Data Size
I Case Study
- I.1 Attribute Value Extraction ()
- I.2 Product Matching ()
- I.3 Product Relation Prediction ()
- I.4 Sentiment Analysis ()
- I.5 Sequential Recommendation ()
- I.6 Multiclass Product Classification ()
- I.7 Product Substitute Identification ()
- I.8 Query Product Ranking ()
- I.9 Answerability Prediction ()
- I.10 Answer Generation ()
@@ Appendix A Task Definition and Data Preprocessing
| Task |
Definition |
Type |
Metrics |
Data |
Product Understanding |
Given the titles, descriptions, features, and brands of the products, extract values for the specific target attributes. |
Information extraction |
precision, recall, F1 (SectionΒ A.1.1) |
MAVEΒ (Yang etΒ al., 2022) based on Amazon Review 2018Β (Ni etΒ al., 2019); OOD: 7 held-out attributes |
Given the titles of two products, predict their relation from ``also buy", ``also view", and ``similar". |
Multi-class classification |
accuracy, macro precision, macro recall, macro F1 |
Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category |
Given the titles, descriptions, manufacturers, and prices of the products from two different platforms, predict if they are the same product. |
Binary classification |
accuracy, precision, recall, F1, specificity, negative prediction rate |
Amazon-Google ProductΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) |
User Understanding |
Given a product review by a user, identify the sentiment that the user expressed on the product. |
Multi-class classification |
accuracy, macro precision, macro recall, macro F1 |
Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category |
Given the interactions of a user over the products, predict the next product that the user would be interested in. |
Ranking |
HR@1 |
Amazon Review 2018Β (Ni etΒ al., 2019) and Amazon Review 2014Β (He & McAuley, 2016; McAuley etΒ al., 2015); OOD: Tools category |
Query Product Matching |
Given a query and a product title, predict the relevance between the query and the product. |
Multi-class classification |
accuracy, macro precision, macro recall, macro F1 |
Shopping Queries DatasetΒ (Reddy etΒ al., 2022) |
Given a user query and a potentially relevant product, predict if the product can serve as a substitute for the userβs query. |
Binary classification |
accuracy, precision, recall, F1, specificity, negative prediction rate |
Shopping Queries DatasetΒ (Reddy etΒ al., 2022) |
Given a user query and a list of potentially relevant products to the query, rank the products according to their relevance to the query. |
Ranking |
NDCG |
Shopping Queries DatasetΒ (Reddy etΒ al., 2022) |
Product QA |
Given a product-related question and reviews of this product, predict if the question is answerable. |
Binary classification |
accuracy, precision, recall, F1, specificity, negative prediction rate |
AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category |
Given a product-related question and reviews as supporting documents, generate the answer to the question. |
Generation |
P, R, F, BLEURT |
AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category |
@@
Raw datasets of the attribute value extraction (, discussed below in Section A.1.1), product matching (, discussed in Section A.1.3), product relation prediction (, discussed in Section A.1.2), and sentiment analysis (, discussed in Section A.2.1) tasks are first split into training, validation, and test data at 8:1:1 ratio. For multi-class product classification (, discussed in Section A.3.1), query product ranking (, discussed in Section A.3.3), product substitute identification (, discussed in Section A.3.2), answerability prediction (, discussed in Section A.4.1), and answer generation (, discussed in Section A.4.2) tasks, the raw datasets are already split. All split data are processed as detailed below and converged with instruction templates into structured data, which will be directly used by LLMs. Based on prior researchΒ (Wei etΒ al., 2021) and considering the high computing demands, we uniformly downsample training sets to a size of 10K, validation sets into 1K, and test sets into 1K size, optimizing data volumes for efficient processing and affordable LLM evaluation.
@@
In , , and tasks, we utilize the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) as the data source, which contains product reviews and metadata in 29 different product categories. For these three tasks, the data from the Electronics, Home, and Sports categories are employed as the IND data sources. We further downsample and combine the processed structured data for each task from these sources at a 1:1:1 ratio into a training, validation, and test set, respectively. For and tasks, we use the data from the AmazonQA datasetΒ (Gupta etΒ al., 2019), which has 17 product categories. To avoid data leakage, we use the data from Sports and Tools categories for the task, and Electronics and Home for task as IND data sources, respectively. Similarly to the previous three tasks, for each of the tasks, we sample and combine the processed data from the IND sources at a 1:1:1 ratio into training, validation, and test sets. For the task, we process all data from Amazon-Google ProductsΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and get a 2K training set, 0.2K validation set, and 0.2K test set. Since this dataset is small, we do not do the downsampling for the task. For , , , and tasks, we directly process and downsample them from the corresponding split in original datasets, separately.
@@
As for OOD datasets, only test data are utilized. The data from the Tools category from the Amazon Review dataset is used as the OOD data source in , , and tasks, while the Cells category from AmazonQA serves as the OOD source in and tasks. For task, we hold out 7 attributes as the OOD dataset. All OOD sources are processed into 1K test sets.
| βTask | Training | Validation | IND Test | OOD Test |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 02,022 | 0,253 | 0,253 | β |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 10,000 | 1,000 | 1,000 | β |
| β | 10,000 | 1,000 | 1,000 | β |
| β | 10,000 | 1,000 | 1,000 | β |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 10,000 | 1,000 | 1,000 | 1,000 |
| β | 92,022 | 9,253 | 10 tasks | 6 tasks |
@@ A.1.1 Attribute Value Extraction ()
where N refers to the total number of entries in ground-truth data. Note that we only consider the prediction result as the correct value (VC) when the ground truth is fully contained in the prediction (e.g. ``bright yellow'' is considered as a correct value with the ground truth of ``yellow''). In our formulation, the precision is calculated as the percentage of correctly predicted positives and negatives in all predictions, and the recall is computed as the percentage of correct positive and negative predictions in all ground truths. We use F1 as the primary metric for the AVE task.
@@
Note that in the above example, for the ``target attribute" ``material", the model should extract its value ``nylon" from the ``product description" and ``product feature".
@@ A.2 User Understanding
The two tasks in this section aim to help the models comprehend the users' needs and preferences.
@@ A.2.2 Sequential Recommendation ()
Given the interactions of a user over the products, predict the next product that the user would be interested in. By learning on this task, the models will have a comprehensive view of user preferences, which enables models to cater to users' future needs.
@@
In the task, we use both product reviews and metadata from the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019). Meanwhile, we use metadata from the Amazon Review 2014 datasetΒ (He & McAuley, 2016; McAuley etΒ al., 2015) as a supplement. The data of Electronics, Home, and Sports categories from both datasets serve as IND sources and the Tools category is used for the OOD test. Moreover, we consider users' review histories as their interactions with products. Following the data processing protocol of UnisRecΒ (Hou etΒ al., 2022), we remove the products without metadata in the 2018 version dataset and conduct the 5-core filter to ensure all products and users appear at least 5 times. After filtering, we sort every user's interactions chronologically and truncate the history with a maximum of 50 products, retaining the least recent history. For the text information of products, we use the metadata from Amazon Review 2014 to fill in the missing information of the same products in the 2018 version metadata.
We also combine the product title, category, and brand to represent a product. The average length of the combined texts is about 21 words. Thus, we retain the first, maximum of 25 words of each combined text for computational efficiency. As in conventional sequential recommendation tasks, we split the last product of the user interactions into the test set as the ground truth next product of the user's interest, the second last product into the validation set, and the remaining products into the training set. When processing the user interactions with instruction templates, for each sample, we randomly select 19 candidate products and mix them up with one ground-truth product. These 20 products serve as the options for the sample.
@@
T: Lorex ACCMIC1 Indoor Audio Microphone Accessory for Surveillance DVR's (Black). Electronics. Camera & Photo. Lorex
@@
Given the interactions shown in the above input, the next product of user's interest for this user is C: Anker iPhone Charger, Powerline Lightning Cable (3ft), MFi Certified for iPhone Xs/XS Max/XR/X
@@ A.3 Query Product Matching
The following three tasks seek to study the relations between the user queries and the potential relevant products to the queries. We use the data from Shopping Queries DatasetΒ (Reddy etΒ al., 2022), which contains 48,300 queries and potentially relevant products to each query. The dataset was originally collected from different market locales in various languages. All products are labeled in four categories: Exact, Substitute, Complement, and Irrelevant, based on their relevance to the query. As in the Shopping Queries Dataset, the four labels are defined as follows:
@@ A.3.1 Multi-class Product Classification ()
Given a query and a product title, predict the relevance between the query and the product (Exact, Substitute, Complement, Irrelevent). This task helps models learn the fine-grained relevance between queries and products, promoting better recommendation results.
@@
We use the product titles to represent products, and translate the Unicode into English. The translated query-product pairs are processed with instruction templates. The ratio of the four labels (Exact, Substitute, Complement, and Irrelevant) in the structured dataset is about 35:10:1:5.
@@
Product: Nike Men's Air Jordan 1 Low White/Gym Red, White/Gym Red/Black, 9
@@ A.4.2 Answer Generation (AG)
Appendix B Instructions
Each task has 6 instructions. One of them is the human-written seed instruction, 4 of them are GPT-4 generated diverse instructions, and the last one is the ``unseen" instruction.
B.1 Instructions on Attribute Value Extraction ()
@@
Generated Instruction 4 Using the product's title, description, features, price, and brand, identify and retrieve the values associated with a specified set of target attributes. Output the extracted values along with their respective sources (e.g., title or feature) indicating where each value was found.
@@
B.2 Instructions on Product Relation Prediction ()

### v2

@@ 6.5 Impact of Training Data Size on
We investigate how the training data size impacts the performance of models. As presented in SectionΒ 3.4, has more than 92 thousand (92K) training samples. In this experiment, we generate 3 smaller training sets of 1K, 10K, and 47K samples. The 1K and 10K samples are generated by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are constructed by randomly selecting 5K samples from each task except for , for which we include all its 2K training samples. The complete results are presented in AppendixΒ I.
@@ Appendix A Task Definition and Data Preprocessing
| Task | Definition | Type | Metrics | Data | |
| Product Understanding | Given the titles, descriptions, features, and brands of the products, extract values for the specific target attributes. | Information extraction | precisionβ, recallβ, F1β (SectionΒ A.1.1) | MAVEΒ (Yang etΒ al., 2022) based on Amazon Review 2018Β (Ni etΒ al., 2019); OOD: 7 held-out attributes | |
| Given the titles of two products, predict their relation from βalso buyβ, βalso viewβ, and βsimilarβ. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category | ||
| Given the titles, descriptions, manufacturers, and prices of the products from two different platforms, predict if they are the same product. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | Amazon-Google ProductΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) | ||
| User Understanding | Given a product review by a user, identify the sentiment that the user expressed on the product. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Amazon Review 2018Β (Ni etΒ al., 2019); OOD: Tools category | |
| Given the interactions of a user over the products, predict the next product that the user would be interested in. | Ranking | HR@1 | Amazon Review 2018Β (Ni etΒ al., 2019) and Amazon Review 2014Β (He & McAuley, 2016; McAuley etΒ al., 2015); OOD: Tools category | ||
| Query Product Matching | Given a query and a product title, predict the relevance between the query and the product. | Multi-class classification | accuracy, macro precision, macro recall, macro F1 | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | |
| Given a user query and a potentially relevant product, predict if the product can serve as a substitute for the userβs query. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | ||
| Given a user query and a list of potentially relevant products to the query, rank the products according to their relevance to the query. | Ranking | NDCG[1] | Shopping Queries DatasetΒ (Reddy etΒ al., 2022) | ||
| Product QA | Given a product-related question and reviews of this product, predict if the question is answerable. | Binary classification | accuracy, precision, recall, F1, specificity, negative prediction rate | AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category | |
| Given a product-related question and reviews as supporting documents, generate the answer to the question. | Generation | P[2], R[2], F[2], BLEURT[3] | AmazonQAΒ (Gupta etΒ al., 2019); OOD: Cells category |
@@
To pursue adherence to data usage requirements, we check the licenses of data sources, ensuring their permission to publish. Table A2 presents the licenses of our curated dataset sources.
| Dataset | License Type | Source |
| Amazon-Google Products | CC-by-4.0 | https://dbs.uni-leipzig.de/research/projects/benchmark-datasets-for-entity-resolution |
| Amazon Review | Not Specified | https://cseweb.ucsd.edu/jΜmcauley/datasets.html#amazon_reviews |
| AmazonQA | Not Specified | https://github.com/amazonqa/amazonqa |
| Shopping Queries Dataset | Apache License 2.0 | https://github.com/amazon-science/esci-data |
@@
Raw datasets of the attribute value extraction (, discussed below in Section A.1.1), product matching (, discussed in Section A.1.3), product relation prediction (, discussed in Section A.1.2), and sentiment analysis (, discussed in Section A.2.1) tasks are first split into training, validation, and test data at 8:1:1 ratio. For multi-class product classification (, discussed in Section A.3.1), query product ranking (, discussed in Section A.3.3), product substitute identification (, discussed in Section A.3.2), answerability prediction (, discussed in Section A.4.1), and answer generation (, discussed in Section A.4.2) tasks, the raw datasets are already split. All split data are processed as detailed below and converged with instruction templates into structured data, which will be directly used by LLMs. Based on prior researchΒ (Wei etΒ al., 2021) and considering the high computing demands, we uniformly downsample training sets to a size of 10K, validation sets to 1K, and test sets to 1K size, optimizing data volumes for efficient processing and affordable LLM evaluation.
@@
In , , and tasks, we utilize the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019) as the data source, which contains product reviews and metadata in 29 different product categories. For these three tasks, the data from the Electronics, Home, and Sports categories are employed as the IND data sources. We further downsample and combine the processed structured data for each task from these sources at a 1:1:1 ratio into a training, validation, and test set, respectively. For and tasks, we use the data from the AmazonQA datasetΒ (Gupta etΒ al., 2019), which has 17 product categories. To avoid data leakage, we use the data from the Sports and Tools categories for the task, and Electronics and Home for task as IND data sources, respectively. Similarly to the previous three tasks, for each of the tasks, we sample and combine the processed data from the IND sources at a 1:1:1 ratio into training, validation, and test sets. For the task, we process all data from Amazon-Google ProductsΒ (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and get a 2K training set, 0.2K validation set, and 0.2K test set. Since this dataset is small, we do not do the downsampling for the task. For , , , and tasks, we directly process and downsample them from the corresponding split in original datasets, separately.
@@
As for OOD datasets, only test data are utilized. The data from the Tools category from the Amazon Review dataset is used as the OOD data source in , , and tasks, while the Cells category from AmazonQA serves as the OOD source in and tasks. For the task, we hold out 7 attributes as the OOD dataset. All OOD sources are processed into 1K test sets.
| Task | Training | Validation | IND Test | OOD Test |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 02,022 | 0,253 | 0,253 | β | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | β | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 10,000 | 1,000 | 1,000 | 1,000 | |
| 92,022 | 9,253 | 10 tasks | 6 tasks |
@@ A.1.1 Attribute Value Extraction ()
where N refers to the total number of entries in ground-truth data. Note that we only consider the prediction result as the correct value (VC) when the ground truth is fully contained in the prediction (e.g. βbright yellowβ is considered as a correct value with the ground truth of βyellowβ). In our formulation, the precisionβ is calculated as the percentage of correctly predicted positives and negatives in all predictions, and the recallβ is computed as the percentage of correct positive and negative predictions in all ground truths. We use F1β as the primary metric for the AVE task.
@@
Note that in the above example, for the βtarget attributeβ βmaterialβ, the model should extract its value βnylonβ from the βproduct descriptionβ and βproduct featureβ.
@@ A.2 User Understanding
The two tasks in this section aim to help the models comprehend the usersβ needs and preferences.
@@ A.2.2 Sequential Recommendation ()
Given the interactions of a user over the products, predict the next product that the user would be interested in. By learning on this task, the models will have a comprehensive view of user preferences, which enables models to cater to usersβ future needs.
@@
In the task, we use both product reviews and metadata from the Amazon Review 2018 datasetΒ (Ni etΒ al., 2019). Meanwhile, we use metadata from the Amazon Review 2014 datasetΒ (He & McAuley, 2016; McAuley etΒ al., 2015) as a supplement. The data of Electronics, Home, and Sports categories from both datasets serve as IND sources and the Tools category is used for the OOD test. Moreover, we consider usersβ review histories as their interactions with products. Following the data processing protocol of UnisRecΒ (Hou etΒ al., 2022), we remove the products without metadata in the 2018 version dataset and conduct the 5-core filter to ensure all products and users appear at least 5 times. After filtering, we sort every userβs interactions chronologically and truncate the history with a maximum of 50 products, retaining the least recent history. For the text information of products, we use the metadata from Amazon Review 2014 to fill in the missing information of the same products in the 2018 version metadata.
We also combine the product title, category, and brand to represent a product. The average length of the combined texts is about 21 words. Thus, we retain the first, maximum of 25 words of each combined text for computational efficiency. As in conventional sequential recommendation tasks, we split the last product of the user interactions into the test set as the ground truth next product of the userβs interest, the second last product into the validation set, and the remaining products into the training set. When processing the user interactions with instruction templates, for each sample, we randomly select 19 candidate products and mix them up with one ground-truth product. These 20 products serve as the options for the sample.
@@
T: Lorex ACCMIC1 Indoor Audio Microphone Accessory for Surveillance DVRβs (Black). Electronics. Camera & Photo. Lorex
@@
Given the interactions shown in the above input, the next product of userβs interest for this user is C: Anker iPhone Charger, Powerline Lightning Cable (3ft), MFi Certified for iPhone Xs/XS Max/XR/X
@@ A.3 Query Product Matching
The following three tasks seek to study the relations between the user queries and the potential relevant products to the queries. We use the data from the Shopping Queries DatasetΒ (Reddy etΒ al., 2022), which contains 48,300 queries and potentially relevant products to each query. The dataset was originally collected from different market locales in various languages. All products are labeled in four categories: Exact, Substitute, Complement, and Irrelevant, based on their relevance to the query. As in the Shopping Queries Dataset, the four labels are defined as follows:
@@ A.3.1 Multi-class Product Classification ()
Given a query and a product title, predict the relevance between the query and the product (Exact, Substitute, Complement, Irrelevant). This task helps models learn the fine-grained relevance between queries and products, promoting better recommendation results.
@@
We use the product titles to represent products and translate the Unicode into English. The translated query-product pairs are processed with instruction templates. The ratio of the four labels (Exact, Substitute, Complement, and Irrelevant) in the structured dataset is about 35:10:1:5.
@@
Product: Nike Menβs Air Jordan 1 Low White/Gym Red, White/Gym Red/Black, 9
@@ A.4.2 Answer Generation (AG)
Appendix B Data Statistics
The comprehensive and wide-ranging tasks in are critical for developing versatile e-commerce foundation models. includes products from 21 categories, such as Sports, Electronics, and Home. This broad coverage ensures that encompasses diverse product information, facilitating the creation of generalist e-commerce models.
Figure A1 shows the distribution of product categories for , , , , , and . The diverse categories in significantly enhance the effectiveness of . Note that the data of , , , and tasks lacks category information, thus their distributions are not shown.
Besides product categories, we also present the distributions of input lengths for each task, measured by word count, in Figure A2. For better clarity, we exclude very long inputs (those representing at most 1% of samples) in the , , , and tasks.
Due to limited computing resources, especially in academic settings, the current dataset has at most 10k training samples per task to fit within the model training capacity. However, scaling up is straightforward by following our data processing procedures detailed in Section 3 and Appendix A. For example, approximately 4.8 million high-quality samples could be further included in the task as outlined in Appendix A.1.1. To allow researchers and developers to customize to their specific needs, all data processing scripts are available at https://github.com/ninglab/eCeLLM/tree/main/data_processing.
Table A4 provides the number of samples at different stages: in the raw data, after data filtering and quality checks, and after sampling. The βAfter Filteringβ column shows the remaining data after applying quality checks. The βAfter Samplingβ column shows the data included in . Detailed statistics for the βAfter Samplingβ column, including training, validation, and testing samples for both in-domain and out-of-domain evaluation, are available in Table A3.
Even after filtering and quality checks, a substantial number of samples remain available for each task, making easily extendable and scalable. We believe can significantly benefit e-commerce foundation model research in various applications.
| Task | Raw Data | After Filtering | After Sampling |
| 15,177,901 | 4,767,579 | 13,000 | |
| 10,868,853 | 7,705,363 | 13,000 | |
| 2,600 | 2,528 | 2,528 | |
| 800,000 | 617,200 | 13,000 | |
| 31,013,687 | 1,946,393 | 13,000 | |
| 2,621,288 | 935,757 | 12,000 | |
| 2,621,288 | 935,757 | 12,000 | |
| 130,652 | 57,707 | 12,000 | |
| 220,541 | 184,814 | 13,000 | |
| 398,654 | 213,612 | 13,000 |
β’
The βAfter Samplingβ column indicates the data included in .
Appendix C Instructions
Each task has 6 instructions. One of them is the human-written seed instruction, 4 of them are GPT-4 generated diverse instructions, and the last one is the βunseenβ instruction.
C.1 Instructions on Attribute Value Extraction ()
@@
Generated Instruction 4 Using the productβs title, description, features, price, and brand, identify and retrieve the values associated with a specified set of target attributes. Output the extracted values along with their respective sources (e.g., title or feature) indicating where each value was found.
@@
C.2 Instructions on Product Relation Prediction ()